# DS communications: Trends, Challenges, and Solutions

### v3

@@ III-D Discussions and Outlook
We stated earlier the different potential modulation, coding, and detection schemes for DS communications. It is worth reiterating that yet experimentation with the efficiency of existing schemes needs to be conducted. DSOC systems are expected to operate at the speed of several tens of gigabits per second and beyond and hence will require a major improvement over existing receiver technology in terms of both data rate and sensitivity [54]. Thanks to their energy efficiency and simplicity, direct-detection systems are preferred over coherent systems. Nevertheless, coherent detection has recently started to get more interest due to its ability to increase the capacity in DS communications [74].
While using single-quadrature homodyne receivers, without a pre-amplifier, resulted in a sensitivity of 1.5 PPB at 156 Mbps [99], demonstrations have shown that pre-amplified erbium-doped fiber coherent receivers can achieve 0.51 PPB increase in sensitivity and 10 Gbps rate [102, 103]. In addition, it is also worth mentioning that there is some focus on in-phase and quadrature modulation homodyne detectors in optical transmission systems for terrestrial as well as DS communications as it promises to achieve higher spectral efficiency than other schemes [104].
@@ IV DSOC Physical Layer Performance Analysis
In order to decide on a modulation type and order, a coding technique, and a detection approach, the channel capacity needs to be determined or at least approximated for a given system setup.
@@ IV-A Capacity Limit
Usually, the Shannon capacity gives a channel maximum error-free data rate in classical information theory [105]. However, in optical communications, channels are best described as quantum channels, shown to accurately describe capacity limit of a corresponding system. For extremely power-limited communications, as it is the case in DS, the capacity limit is best characterised by Holevo theorem, which approximates the solution of the quantum channels capacity limit [106, 107]. Based on Holevo theorem [85], the analytical closed-form expression of the quantum capacity, which asymptotically expresses the dimensional information efficiency, measured in bits/dimension, is given by
@@
where $c_{p}$ is photon information efficiency and e is the exponential constant. In theory, the optimal coherent-state modulation achieves the ultimate quantum limit, which is the capacity limit that can possibly be approached by a system based on Holevo limit approximation. Still, practical sub-optimal coherent receivers are operating in near the ultimate Holevo limit but in the low photon efficiency region as they hit a brick wall of $1.44$ bits/photon and hence they are not efficient for DS applications [85].
@@
where $C_{PCR}$ is given by (3). Dolinar et al. in [108] proposed using coherent states with PC to get closer to the Holevo limit; thanks to its ability to produce coherent states. This is challenging and not practical from an implementation perspective, though. Quantum number state (NS), also known as Fock state [109], communications could be an alternative, where the quantum state is with a well-defined number of photons are shown to be able to reach the Holevo limit [108]. This could be challenging in practice, again though, since it requires high channel transmissivity, which is the probability of transmitted NS photon being received at the detector. To further illustrate this, considering OOK modulation, the single-photon NS capacity at high photon information efficiency can be written in terms of the channel transmissivity, denoted by $\eta$, asymptotically as
@@
As shown in Fig. 5, PPM cannot reach the Holevo limit even when $\eta=1$. With a transmissivity of one, NS along with OOK gets to the ultimate limit. As the quantum NS transmissivity gets higher, the capacity gets closer to the Holevo limit in the case of OOK. We also show in the same figure the capacity limit for PPM-PC, and quantum NS modulation with OOK and PPM.
@@
As a way to analyze the DSOC links performance for a PPM-PC detector, in Table III, we show the soft-decision capacity at various planets average distance. Soft-decision capacity corresponds to implementing a soft-decision receiver which passes on slot counts and use them to determine the probability of sending a certain sequence of bits [54]. Here we assume that the transmitter and the receiver's aperture diameters to be 0.22 m and 4 m, respectively, a transmitted power of 4 W [28, 38, 68], a 0.25 ns time slot, a PPM modulation order of 16, and a noise power of 1.1620$\times$10-16 W. The losses of Earth to Mars communications, for example in the considered setup, are as follows: a path-loss of  366 dB, the sum of other losses for the received power not including the detection loss is 6.6 dB, and the detection loss is 4.35 dB. We notice that the capacity dramatically gets smaller as we get to further planet distances.
@@ IV-B Discussions and Outlooks
Although direct detectors can not reach capacity limit, the capacity can be improved by increasing the detector sensitivity, selecting a modulation format with a high peak to average power ratio like PPM, and reducing the amount of background light detected. Compact array detectors, which decrease blocking and pointing losses and help with beam acquisition and tracking process [110], achieve higher data rates than a single detector, only for limited distance though [28]. In addition, while being flexible and scalable, multiple receivers are considered more cost-effective than single receivers for data rate improvement [92]. However, the single receiver could outperform the array in case of no AT [111].
@@ V DSOC Networking Protocols
In addition to the different physical layer techniques and their impact on DSOC performance as we discussed in the sections III and IV, networking protocols are of paramount importance to make the communications links reliable and overcome a number of different complex challenges facing DSOC discussed in Section II. Seamless coordination between satellites, ground communications systems, and DS probes is needed. Matured terrestrial network protocols are mostly inefficient in DS environments where distances are extreme and connections are intermittent; hence, as suggested in [112], calling for well-tailored network protocols is essential. NASA, for example, relies on delay/disruption tolerant networking (DTN) [50, 113, 114] architecture as an intuitive solution for space missions. The architecture can handle excessive delays and fast-changing environment while maintaining reliability.
The Consultative Committee for Space Data Systems (CCSDS), which is one of the most important organizations that discuss and set standards for space data systems, recommends architectures for space networks, such as space communications protocol specifications (SCPS) architecture [115], space packet protocol (SP)-based architecture [116], and IP architecture. Due to its ability to handle long and variable delays and link intermittent connectivity issues, DTN, however, is the most persuasive architecture. The automatic store-and-forward feature of DTN, which allows packets to be forwarded to the next destination only if the availability of the corresponding link is ensured, is making the DTN architecture a natural choice for severe environments [117, 118].
@@ V-A DTN Architecture Layers and Protocols
Application layer: In this layer, CCSDS file delivery protocol (CFDP) is a widely used protocol [119]. In CFDP, the data units are delivered independently from each other, aiming to increase the amount of communicating data [120, 121]. Space missions, however, do not only rely on file transfer for all their communications requirements. CCSDS proposed messaging exchange protocol called asynchronous message service (AMS), which is suitable, for example, in real-time control and robotic operation collaboration. AMS is designed to reduce the communications burden of space missions and hence their cost and risk of failure [122].
Bundle layer: Next to the application layer, there is the bundle layer. This layer distinguishes DTN from the conventional TCP/IP architecture. Due to frequent disruptions with long transmission delays, in a hop-by-hop transport fashion, the bundle layer that only has the bundle protocol (BP) adds a store-and-forward mechanism for each DTN node, i.e., every engine that runs BP in the network, and sends data in large packets called bundles. The BP mechanism is different than end-to-end protocols such as TCP, where reliability is achieved through end-to-end acknowledgments and retransmission excluding the intermediate nodes [123]. TCP or end-to-end protocols have a low transport efficiency if used in the DS network [50]. The hop-by-hop principle, however, enhances transport efficiency in such challenging environments [50]. To enhance reliability and reduce congestion, BP supports custody transfer which is an automatic retransmission request (ARQ) mechanism that takes custody of a bundle, node by node, until it reaches its destination [124]. ARQ mechanisms are implemented in the data link layer protocol as they are commonly used, however, in addition to that, they are implemented in the upper layers of the DTN network architecture [123], to ensure reliability of data delivery. An optional end-to-end acknowledgment feature [125] could be used for situations where original data senders are hard to reach for a request of retransmission [126]. Furthermore, bundles are carried out through convergence layer adapters (CLAs) which allows DTN to encompose different underlying protocols and hence work well in a heterogeneous DS networks [127]. Conventional routing protocols, most built for the TCP/IP architecture, are not efficient for time-variant DS networks [50]. In efforts to tackle these challenges, contact graph routing (CGR) protocol has been developed for space networks based on DTN architecture [128]. CGR exploits the mission communications operation plans to construct a contact graph that maps time-varying network connections which is used as a consequence for bundles' best route calculation.
@@
Layers
Protocol
Functions & Mechanism
Advantages
Disadvantages
Application
CFDP
It is capable of file transfer, delivery, and storage management services, namely loading, dumping, and controlling [129].
It assumes that communications entities have storage mediums called filestore. The protocol operates by copying data between these mediums. It has a delivery process that gives the capability of transferring files in arbitrary networks and works in a reliable as well as an unreliable mode
Scalable, minimizes the resources required to only the necessary elements for file transfer operation [129]
Causes overhead in the network when used along with some other subsequent protocols [41, 119], does not preform well for one-directional transmission long-delay communications link [130]
AMS
AMS provides a asynchronous message exchange services [131]. It is designed to make the functioning parts of an application, called modules, operate in isolation and make communications between these modules self-configuring. Hence complexity of the modular data system is reduced
Simple to use, highly automated, flexible, robust, scalable, efficient [122]
Bundle
Bundle
It applies overlay network principle which provides hop-by-hop transmission for data delivery and status reporting [124]. Information is sent in bundles, large groups of data, in a store and forward fashion with optional custody transfer [132]
Data accountability, reliability in heterogeneous networks [124]
Relatively high memory consumption [124], not 100% reliable [133]
Transport
LTP
It provides optional reliable communications, as it divides blocks of data into red-part and green-part where the former need to be delivered reliably but unnecessarily the latter [134]
High throughput for the application level data, i.e. high good- put [135]
In-order delivery is not supported, low performance for short distance and small delay, with low BER link [135]
Internet
IPoC
IPoC specifies the implementation of the IP protocol over CCSDS SDLPs. The protocol data units (PDU) of IP get encapsulated by EP and then sent for framing by SDLP [136]
IPv4 and IPv6 are commercially available technologies directly implemented in space, with good transfer performance and cost-effectiveness for short links [123]
Routing fails in long transmission delays [136]
Data link
USLP
It utilizes features and functions of existing CCSDS data link protocols which are Prox-1, TC, TM, and AOS, with added advances. It was developed to meet the requirements of space missions for efficient transfer of data [137]
Designed to reduce cost and complexity, increases data rate, and meet the required data rate expected for DSOC, which can not be reached with the previous CCSDS standards [137]
Physical
CCSDS-HPE
It specifies the characteristics of the transmitted laser signal [138]
@@
Internet layer: When it comes to addressing nodes and packet delivery, i.e., internet layer functionalities, nodes that are communicating in the same region, protocols like IP can be used [145]. In an attempt to make IP more compatible with the space environment, however, IP over CCSDS space links (IPoC) is recommended [136]. If a sender is located in a region different than the destination, however, e.g. two different planets, bundles are addressed differently for more effective routing. Addressing schemes like IP in general indicate the geographic location of the host. Such methods, however, provide no information about the host itself. In the case of interplanetary communications, the location of the host might change during the course of delivering the data, and hence information about the host itself is beneficial for accurately assigning destinations [146]. A naming structure called universal resource identifiers (URI), also referred to as endpoint identifiers (EID), where each DTN node gets associated with an identifier, a type of address that specifies the region, application, and host [127], plays an important rule for achieving efficient interplanetary communications.
@@
Data link layer: Packets from the internet layer get encapsulated either by space packets (SP[116] or encapsulation packets (EP) protocols [147]. Afterwards, packets go through the framing process in the data link layer with protocols provided by the space data link protocols (SDLPs). These protocols are telecommand (TC[148], telemetry (TM) [149], proximity-1 (Prox-1) [150], advanced orbiting systems (AOS) [151], and unified space data link protocol (USLP) [152]. CCSDS SDLPs are responsible for transferring data with different types and variable lengths in space links. TM, AOS, and USLP can be used for optical communications [138]. Frames go through coding and synchronization data link sub-layer to prepare them to the physical layer, which transmits them to the medium based on the physical layer technical standard.
@@
Physical layer: For DSOC, high photon efficiency (HPE) specifications are recommended by CCSDS to be used in the physical layer. This standard characterizes the transmitted laser signal, including used modulation and related parameters, for space-ground and ground-space links [138].
In Table IV, we present some other aspects of the most significant and advanced DS protocols discussed above corresponding to their DTN layers.
Figure 6: BP custody transfer, CLAs, and CGR dynamics visualization
@@
To help visualize how DS DTN defining features work, Fig. (6) further explains the concept of CLA, CGR, EIDs, and custody transfer. Let us build a scenario for transmitting information from a mission control node located on Earth N1 to a Mars rover N5. To have that happen, a bundle, denoted by B, is created with the destination EID that includes destination region and host in addition to the underlying application, App. A contact plan is also created giving an estimation of future connection to nodes based on their locations and characteristics. The contact plan can be kept at N1 along with the routing table for a centralized routing [153]. The CGR algorithm at N1 uses the contact plan to estimate possible routes by evaluating parameters such as local time, delivery time, bundle best transmission time, bundle priority, and bundle queue status. In this scenario, it is assumed that the best route to reach N5 is through node N2, N3, and then N4. As the bundle B travels over networks, it gets handled with CLAs to integrate different transport and other underlying protocols if any. At N3, B checks N4 which is assumed to be unavailable at first contact, time t1. As a consequence, B is held in custody until the node becomes, at time t2, accessible and hence B finds its way to reach the destination.

### v4

@@ III-D Discussions and Outlook
We stated earlier the different potential modulation, coding, and detection schemes for DS communications. It is worth reiterating that yet experimenting with existing schemes efficiency needs to be conducted. DSOC systems are expected to operate at the speed of several tens of gigabits per second and beyond and hence will require a major improvement over existing receiver technology in terms of both data rate and sensitivity [54]. Thanks to their energy efficiency and simplicity, direct detection systems are preferred over coherent systems. Nevertheless, coherent detection has recently started to get more interest due to its ability to increase the capacity of DS communications [75].
While using SQ homodyne receivers, without a pre-amplifier, resulted in a sensitivity of 1.5 PPB at 156 Mbps [102], demonstrations have shown that pre-amplified erbium-doped fiber coherent receivers can achieve a 0.51 PPB increase in sensitivity and a 10 Gbps rate [105, 106]. In addition, it is also worth mentioning that there is some focus on in-phase and quadrature modulation homodyne detectors in optical transmission systems for terrestrial as well as DS communications as they promise to achieve higher-spectral efficiency than other schemes [107].
@@ IV DSOC Physical Layer Performance Analysis
To decide on a modulation type, coding technique, and detection approach, the channel capacity needs to be determined or at least approximated for a given system setup.
@@ IV-A Capacity Limit
Shannon capacity gives a channel the maximum error free data rate in classical information theory [108]. However, in optical communications, channels are best described as quantum channels, which are shown to accurately describe capacity limit of a corresponding system. For extremely power-limited communications, as is the case in DS, the capacity limit is best characterized by the Holevo theorem, which approximates the solution of the quantum channel capacity limit [109, 110]. Based on the Holevo theorem [86], the analytical closed-form expression of the quantum capacity, which asymptotically expresses the dimensional information efficiency, measured in bits/dimension, is given by
@@
where $c_{p}$ is the photon information efficiency (PIE) which is the amount of information that can be transmitted per photon, measured in bits/photon, and $e$ is the exponential constant. In theory, the optimal coherent state modulation achieves the ultimate quantum limit, which is the capacity limit that can possibly be approached by a system based on the Holevo limit approximation. Still, practical sub-optimal coherent receivers are operating near the ultimate Holevo limit but in the low photon efficiency region as they hit a brick wall of $1.44$ bits/photon and hence they are not efficient for DS applications [86].
@@
where $C_{PCR}$ is given by (3). Dolinar et al. [111] proposed using coherent states with PC to get closer to the Holevo limit. This is challenging and not practical from an implementation perspective, though. Quantum number state (NS), also known as Fock state [112], communications could be an alternative, where the quantum state with a well-defined number of photons are shown to be able to reach the Holevo limit [111]. This could be challenging in practice, again, since it requires high-channel transmissivity, which is the probability of transmitted NS photon being received at the detector. To further illustrate this, considering OOK modulation, the single photon NS capacity at high PIE can be written in terms of the channel transmissivity, denoted by $\eta$, asymptotically as
@@
As shown in Fig. 5, PPM cannot reach the Holevo limit even when $\eta=1$. With a transmissivity of one, NS along with OOK gets to the ultimate limit. As the quantum NS transmissivity gets higher, the capacity gets closer to the Holevo limit in the case of OOK. We also show in the same figure the capacity limit for PPM-PC, and quantum NS modulation, with OOK and PPM.
@@
To analyze the DSOC link performance for a PPM-PC detector, in Table III, we show the soft decision capacity at various planets average distances. Soft decision capacity corresponds to implementing a soft decision receiver that passes on slot counts and uses them to determine the probability of sending a certain sequence of bits [54]. Here we assume that the transmitter and the receiver aperture diameters are 0.22 m and 4 m, respectively, with a transmitted power of 4 W [28, 38, 69], 0.25 ns time slot, PPM modulation order of 16, and noise power of 1.1620$\times$10-16 W. The losses of Earth to Mars communications, for example in the considered setup, are as follows: a path loss of  366 dB, the sum of other losses for the received power not including the detection loss is 6.6 dB, and the detection loss is 4.35 dB. We notice that the capacity dramatically gets smaller as we get to further planet distances.
@@ IV-B Discussions and Outlooks
Although direct detectors cannot reach capacity limit, the capacity can be improved by other techniques including increasing sensitivity and reducing the amount of background light detected. All relevant emerging technologies would be of interest as they can vitally define the future capabilities in DS environment. One of the innovative solutions that emerges to improve capacity limit and data rate, flexibility, scalability, and cost-effectiveness is compact array detectors [113, 28]. Array detectors decrease blocking and pointing losses and help with beam acquisition and tracking processes. It has been shown that, with the use of SNSPD, arranging a couple to tens of detectors significantly reduces the blocking loss, increases efficiency to 93% [114] with up to 7.41 PPB sensitivity [115], and utilizes data rates in the range of gigabits per second [116]. Such detectors hold a continuous active area, ready to receive the upcoming photons, which could be as low as 3 $\mu$m by 3.3 $\mu$[116]. Although smaller active areas result in higher-data rates, increasing them improves system detection efficiency. Hence, there seem to be efforts, see e.g. [117, 115], to get the benefits of relatively large-active areas without sacrificing the data rate and facing the reading out challenge which results in increasing temperature and complexity. Despite the promises of array detectors, according to [118, 28], in some cases the single receiver could outperform the array, for relatively long distances and in case of no AT.
There are a number of other technologies suggested regarding increasing the effectiveness of noise rejection, hence, improving PIE upper limit. That includes the use of nonlinear optical quantum pulse gating (QPG) technique [119], which can filter out the unwanted temporal noise. Even more improvements have been proven through the utilization of photon number resolved (PNR) detectors which interestingly, in contrast to conventional PCs for instance individual SNSPDs [120], are able to read more than one photon simultaneously. PNR detectors are highly regarded for quantum optic applications and quantum communications. Making this type of detector along with other physical layer optimal solutions to yet be investigated for further improvements in terms of capacity limit.
@@ V DSOC Networking Protocols
In addition to the different physical layer techniques and their impact on DSOC performance as we discussed in the sections III and IV, networking protocols are of paramount importance to make the communications links reliable and overcome a number of different complex challenges facing DSOC discussed in Section II. Seamless coordination between satellites, ground communications systems, and DS probes is needed. Matured terrestrial network protocols are mostly inefficient in DS environments where distances are extreme and connections are intermittent; hence, as suggested in [121], calling for well-tailored network protocols is essential. NASA, for example, relies on delay/disruption tolerant networking (DTN) architecture [50, 122, 123] as an intuitive solution for space missions. The architecture can handle excessive delays and fast-changing environment while maintaining reliability.
The Consultative Committee for Space Data Systems (CCSDS), which is one of the most important organizations that discuss and set standards for space data systems, recommends architectures for space networks including DTN, space communications protocol specifications (SCPS) architecture [124], space packet protocol (SP)-based architecture [125], and IP architecture. Rendering to its ability to handle long and variable delays and link intermittent connectivity issues, DTN is the most persuasive architecture. The automatic store-and-forward feature of DTN, which allows packets to be forwarded to the next destination only if the availability of the corresponding link is ensured, is making the DTN architecture a natural choice for severe environments [126, 127].
@@ V-A DTN Architecture Layers and Protocols
Application layer: In this layer, CCSDS file delivery protocol (CFDP) is a widely used protocol [128]. In CFDP, the data units are delivered independently from each other, aiming to increase the amount of data [129, 130]. Space missions, however, do not only rely on file transfers for all their communications requirements. CCSDS proposed a messaging exchange protocol called asynchronous message service (AMS) suitable for various applications including real time control and robotic operation collaboration. AMS is designed to reduce the communications burden of space missions and hence their cost and risk of failure [131].
Bundle layer: Next to the application layer, there is the bundle layer. This layer distinguishes DTN from the conventional TCP/IP architecture. Due to frequent disruptions with long transmission delays, in a hop-by-hop transport fashion, the bundle layer that only has the bundle protocol (BP) adds a store-and-forward mechanism for each DTN node, i.e., every engine that runs BP in the network, and sends data in large packets called bundles. The BP mechanism is different from end-to-end protocols such as TCP, where reliability is achieved through end-to-end acknowledgments and retransmission excluding the intermediate nodes [132]. TCP or end-to-end protocols have a low transport efficiency if used in the DS network; the hop-by-hop principle, however, enhances transport efficiency in such challenging environments [50]. To enhance reliability and reduce congestion, BP supports custody transfer which is an automatic retransmission request (ARQ) mechanism that takes custody of a bundle, node-by-node, until it reaches its destination [133]. ARQ mechanisms are implemented in the data link layer protocol as they are commonly used; however, in addition to that, they are implemented in the upper layers of the DTN network architecture [132], to ensure the reliability of data delivery. An optional end-to-end acknowledgment feature [134] could be used in situations where original data senders are hard to reach for a request of retransmission [135]. Furthermore, bundles are carried out through convergence layer adapters (CLAs) which allow DTN to encompass different underlying protocols and hence work well in heterogeneous DS networks [136]. Conventional routing protocols, most built for the TCP/IP architecture, are not efficient for time-variant DS networks [50]. In efforts to tackle these challenges, contact graph routing (CGR) protocol has been developed for space networks based on DTN architecture [137]. CGR exploits the mission communications operation plans to construct a contact graph that maps time-varying network connections which is used as a consequence for bundles’ best route calculation.
@@
Layers Protocol Functions & Mechanism Advantages Disadvantages
Application CFDP It is capable of file transfer, delivery, and storage management services, namely loading, dumping, and controlling [138]. It assumes that communications entities have storage mediums called filestore. The protocol operates by copying data between these mediums. It has a delivery process that gives the capability of transferring files in arbitrary networks and works in a reliable as well as an unreliable mode Scalable, minimizes the resources required to only the necessary elements for file transfer operation [138] Causes overhead in the network when used along with some other subsequent protocols [41, 128], does not perform well for one-directional transmission long delay communications link [139]
AMS AMS provides a asynchronous message exchange services [140]. It is designed to make the functioning parts of an application, called modules, operate in isolation and make communications between these modules self configuring. Hence complexity of the modular data system is reduced Simple to use, highly automated, flexible, robust, scalable, efficient [131] -
Bundle Bundle It applies overlay network principle which provides hop-by-hop transmission for data delivery and status reporting [133]. Information is sent in bundles, large groups of data, in a store-and-forward fashion with optional custody transfer [141] Data accountability, reliability in heterogeneous networks [133] Relatively high-memory consumption [133], not 100% reliable [142]
Transport LTP It provides optional reliable communications, as it divides blocks of data into red-part and green-part where the former needs to be delivered reliably but unnecessarily the latter [143] High throughput for the application level data, i.e., high good-put [144] In-order delivery is not supported, low performance for short distance and delay, with low BER link [144]
Internet IPoC IPoC specifies the implementation of the IP protocol over CCSDS SDLPs. The protocol data units (PDU) of IP get encapsulated by EP and then sent for framing by SDLP [145] IPv4 and IPv6 are commercially available technologies directly implemented in space, with good transfer performance and cost effectiveness for short links [132] Routing protocols fail in links with long delay that is caused by the very long distance [145]
Data link USLP It utilizes features and functions of existing CCSDS data link protocols which are Prox-1, TC, TM, and AOS, with added advances. It was developed to meet the requirements of space missions for efficient transfer of data [146] Designed to reduce cost and complexity, increases data rate, and meet the required data rate expected for DSOC, which can not be reached with the previous CCSDS standards [146] -
Physical CCSDS-HPE It specifies the characteristics of the transmitted laser signal [147] - -
@@
Internet layer: When it comes to addressing nodes and packet delivery, i.e., internet layer functionalities, nodes that are communicating in the same region, protocols like IP can be used [154]. In an attempt to make IP more compatible with the space environment, however, IP over CCSDS space links (IPoC) is recommended [145]. If a sender is located in a region different from the destination, however, e.g. two different planets, bundles are addressed differently for more effective routing. Addressing schemes like IP in general indicate the geographic location of the host. Such methods, however, provide no information about the host itself. In the case of interplanetary communications, the location of the host might change during the course of delivering the data, and hence information about the host itself is beneficial for accurately assigning destinations [155]. A naming structure called universal resource identifiers (URI), also referred to as endpoint identifiers (EID), where each DTN node gets associated with an identifier, a type of address that specifies the region, application, and host [136], plays an important role in achieving efficient interplanetary communications.
@@
Data link layer: Packets from the internet layer get encapsulated either by space packets (SP[125] or encapsulation packets (EP) protocols [156]. Afterwards, packets go through the framing process in the data link layer with protocols provided by the space data link protocols (SDLPs). These protocols are telecommand (TC[157], telemetry (TM) [158], proximity-1 (Prox-1) [159], advanced orbiting systems (AOS[160], and unified space data link protocol (USLP) [161]. CCSDS SDLPs are responsible for transferring data with different types and variable lengths in space links. TM, AOS, and USLP can be used for optical communications [147]. Frames go through coding and synchronization data link sub-layer to prepare them for the physical layer, which transmits them to the medium based on the technical standard of the physical layer .
@@
Physical layer: For DSOC, high-photon efficiency (HPE) specifications are recommended by CCSDS to be used in the physical layer. This standard characterizes the transmitted laser signal, including modulation and related parameters, for space-ground and ground-space links [147].
Figure 6: BP custody transfer, CLAs, and CGR dynamics visualization.
@@
In Table IV, we present some other aspects of the most significant and advanced DS protocols discussed above corresponding to their DTN layers. Additionally, to help visualize how DS DTN features work, Fig. 6 further explains the concepts of CLA, CGR, EID, and custody transfer. Let us build a scenario for transmitting information from a mission control node located on Earth N1 to a Mars rover N5. To have that happen, a bundle, denoted by B, is created with the destination EID that includes the destination region and host in addition to the underlying application, App. A contact plan is also created giving an estimation of future connections to nodes based on their locations and characteristics. The contact plan can be kept at N1 along with the routing table for centralized routing [162]. The CGR algorithm at N1 uses the contact plan to estimate possible routes by evaluating parameters such as local time, delivery time, bundle best transmission time, bundle priority, and bundle queue status. In this scenario, it is assumed that the best route to reach N5 is through nodes N2, N3, and then N4. As bundle B travels over networks, it gets handled with CLAs to integrate different transport and other underlying protocols, if any. At N3, B checks N4 which is assumed to be unavailable at first contact, time t1. As a consequence, B is held in custody until the node becomes, at time t2, accessible and hence B finds its way to reach the destination.